# Closed mean curvature flows with asymptotically conical singularities

### v2

@@ 6. Ważewski’s box argument
Angenent–Ilmanen–Velázquez constructed $n$-dimensional asymptotically conical shrinkers such that level set flows starting from their asymptotic cones fatten for $n\in\left\{3,4,5,6\right\}$ in [AIV, Theorem 1]. [^1] The proof of Theorem 1 in [AIV] was provided in Section 3 (for proving Theorem 5) and Section 5.1 there. Take such a shrinker $\Sigma$ and let $\mathcal{C}$ be its asymptotic cone. Theorem 1 implies that there exists a smooth closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the level set flow $M_{t}$ starting from $M$ develops a singularity at $({\bf 0},0)\in\mathbb{R}^{n+1}\times\mathbb{R}$ whose tangent flow is modeled on $\Sigma.$ Thus, $M_{0}$ is a hypersurface with a singularity at the origin modeled on the cone $C$ (cf. [CS, Corollary 1.2]). Since the level set flow starting from $\mathcal{C}$ fattens, the main result in [CDS] implies that the level set flow $M_{t}$ fattens for positive time. ∎

### v3

@@ 6. Ważewski’s box argument
Angenent–Ilmanen–Velázquez constructed $n$-dimensional asymptotically conical shrinkers such that level set flows starting from their asymptotic cones fatten for $n\in\left\{3,4,5,6\right\}$ in [AIV, Theorem 1]. [^2] The proof of Theorem 1 in [AIV] was provided in Section 3 (for proving Theorem 5) and Section 5.1 there. Take such a shrinker $\Sigma$ and let $\mathcal{C}$ be its asymptotic cone. Theorem 1 implies that there exists a smooth closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the level set flow $M_{t}$ starting from $M$ develops a singularity at $({\bf 0},0)\in\mathbb{R}^{n+1}\times\mathbb{R}$ whose tangent flow is modeled on $\Sigma.$ Thus, $M_{0}$ is a hypersurface with a singularity at the origin modeled on the cone $\mathcal{C}$ (cf. [CS, Corollary 1.2]). Since the level set flow starting from $\mathcal{C}$ fattens, the main result in [CDS] implies that the level set flow $M_{t}$ fattens for positive time. ∎